Find the probability of the following. A bag contains 6 marbles, 9 blue marbles, and 4 green marbles. What is the probability of that a randomly selected marble is green? Please solve and show work.
Thank you very much!

Answer by Cromlix(4381)

Number of green marbles = 4
Total number of marbles = 6+9+4 = 19
Therefore the probability of selecting a green marble
= 4/19
4 in 19 chance.
Hope this helps:-)